# Hierarchical certification of non-classical network correlations

### v3

@@ 6 Inequalities for general networks
$\sum_{\vec{t}}M^{\vec{t}}\mathcal{S}^{(\vec{t})}(B_{x_{i_{1}}}^{(\vec{t})},\dots,B_{x_{i_{k}}}^{(\vec{t})})\leq 2^{k-1}.$ (29)
7 Noise robustness
@@
$\rho_{v_{i}}=v_{i}|\phi_{i}\rangle\langle{\phi_{i}}|+\frac{1-v_{i}}{d^{2}}I_{d^{2}},$ (30)
@@
Using the states $|\phi_{i}\rangle$ and the measurements that achieve the quantum violations described throughout the manuscript (see, e.g., Section 3 and Appendix A.3), the corresponding quantities $I$ and $J$ get multiplied by a factor $\prod_{i=1}^{m}v_{i}$. This implies that, in order to observe a violation, the visibilities of the sources must satisfy $\prod_{i=1}^{m}v_{i}>\frac{1}{\sqrt{2}}$ when using maximally entangled states. Similarly, using maximally entangled states for $|\phi_{i}\rangle$ in the $n+1$-partite star network, $\ell$-NN correlations can be detected via the violation of Eq. (23) for $\prod_{i=1}^{m}v_{i}>2^{-\left(\ell-\frac{n}{2}\right)}$, while $\ell$-QNN correlations can be detected via the violation of Eq. (20) whenever $\prod_{i=1}^{m}v_{i}>2^{-\frac{\ell}{2}}$. As discussed in Section 3, this shows that Eq. (23) cannot be used for detecting FNN (although it can be used for detecting $\ell$-NN as long as $\ell>n/2$), while Eq. (20) can be used for detecting $\ell$-QNN in a noise-robust manner independently of the number of sources in the network, since the critical visibility only depends on $\ell$.
@@
$\prod_{i=1}^{n}v_{i}>\frac{1}{\sqrt{1+\prod_{j=1}^{n-1}\sin^{2}2\theta_{j}}}.$ (31)
Analogously, when decomposing star networks into sets of of tripartite chain networks, FNN can in principle be guaranteed when every pair of visibilities satisfies
$v_{i}v_{j}>\frac{1}{\sqrt{1+\sin^{2}2\theta_{i}\sin^{2}2\theta_{j}}}.$ (32)
@@
8 Discussion
In this work we have provided Bell-like inequalities that witness genuinely quantum phenomena in networks, in the sense that their violations by particular distributions guarantee that these cannot be reproduced when at least a number of the sources in the network are classical. We have done so using two different assumptions on the non-classical sources, namely (i) that they are only constrained by the no-signaling principle, and (ii) that they distribute systems described by quantum mechanics. The former leads to a hierarchy of levels of network nonlocality, that interpolates from the standard definition of network nonlocality [23] to that of full network nonlocality [35]. The latter gives rise to an analogous hierarchy, that has looser requirements for experimental observations and is thus more suitable for applications in quantum communication. Concretely, we have provided Bell-like inequalities whose violation guarantees (i) $n/2$-QNN and $\ell$-NN in the $n$-partite chain network (Eq. (17), and Eq. (40) in Appendix A, respectively), (ii) $\ell$-NN and $\ell$-QNN for any $\ell$ in the $n+1$-partite star network (Eqs. (23) and (20), respectively), and (iii) FNN in arbitrary chain and star networks (Eqs. (26) and (28), respectively).
Being able to guarantee, from observed statistics, the absence of classical sources in a network is an important step in the development of network-based quantum information protocols. Concrete protocols exist, for instance for quantum conference key agreement [49, 50, 51] or quantum secret sharing [52, 53]. However, their proofs of security rely on guaranteeing global multipartite non-classicality, which leads to demanding requirements. Our work paves the way to the development of protocols and security proofs that exploit the network structure, which has received limited attention so far [54, 55, 53]. Indeed, consider the $n+1$-partite star network. Following Refs. [54, 53] one can bound the predictive power of an adversary for all outcomes $\vec{a}$ by the value obtained in Eq. (20) via
$D(P(\vec{e}|\vec{a};\vec{x},y,\vec{z}),\prod_{i=1}^{n}P(e_{i}|z_{i}))\leq(2-\mathcal{S}_{\ell})n,$ (33)
@@ 8 Discussion
Having inequalities for detecting $\ell$-NN and $\ell$-QNN allows to easily see how inserting the assumption that nature is governed by quantum mechanics results in milder conditions to guarantee non-classicality, which is relevant for experimental realizations. For instance, detection of FNN ($\ell=1$) in the star network with $n=3$ via Eq. (23) is not possible with noisy versions of the generalized EPR states considered throughout for any visibility, while detection of FQNN is possible when the product of the states’ visibilities exceeds $2^{-1/6}\approx 0.891$.
@@
Acknowledgements
This work was supported by the National Natural Science Foundation of China (Nos. 62172341, 61772437), Sichuan Natural Science Foundation (No. 2023NSFSC0447), Interdisciplinary Research of Southwest Jiaotong University China (No.2682022KJ004), the NCCR SwissMAP, the Spanish Ministry of Science and Innovation MCIN/AEI/10.13039/ 501100011033 (CEX2019-000904-S and PID2020-113523GB-I00), the Spanish Ministry of Economic Affairs and Digital Transformation (project QUANTUM ENIA, as part of the Recovery, Transformation and Resilience Plan, funded by EU program NextGenerationEU), Comunidad de Madrid (QUITEMAD-CM P2018/TCS-4342), Universidad Complutense de Madrid (FEI-EU-22-06), and the CSIC Quantum Technologies Platform PTI-001.
@@ Appendix A Nonlinear inequalities for chain networks
In this appendix we analyze the cases that are missing in Section 2.2, namely, $\ell$-QNN in chain networks with an even number of parties, and $\ell$-NN in chain networks with an arbitrary number of parties. The core of the derivations is based on the concept of $k$-independent sets of sources [31].
@@ A.1 $\ell$-QNL bounds for even number of parties
Let us begin discussing the case when the non-classical sources distribute quantum systems. An inequality for the case of chains with an odd number of parties was derived in Section 2.2, namely
$I^{\frac{2}{n+1}}+J^{\frac{2}{n+1}}\leq\begin{cases}2^{\frac{n-1}{2(n+1)}}&\ell>\frac{n}{2}\\ \sqrt{2}&\ell<\frac{n}{2}\end{cases}.$ (34)
@@
If less than half of the sources in the network are classical, the situation is similar to that in Eq. (16). Namely, the joint correlations generated by either $\mathcal{N}_{1}$ or $\mathcal{N}_{2}$ achieve the maximal bound of $\sqrt{2}$ for the inequality (15) with $n-1$ parties, i.e., $I^{\frac{2}{n}}+J^{\frac{2}{n}}\leq\sqrt{2}$. This means that, in the case of an even number of parties and less than half of the sources being classical, it is not currently possible to witness $\ell$-QNN with $\ell<n/2$.
@@ A.2 $\ell$-NN bounds
Now we focus on the bounds that are achieved when the only constraint on the non-classical sources is that they distribute systems that satisfy the non-signaling principle. As in the previous section, we distinguish the cases networks with odd and even number of parties.
@@
Suppose that the chain network consists of $\ell$ classical variables with $\ell\geq 1$ while the other sources distribute arbitrary no-signaling systems, as shown in Fig. 2(b). Equation (11) holds for this kind of networks. However, in contrast with the derivation in Section 2.2, for sources distributing non-signaling systems the objects in Eq. (13) are bounded by
$\left|\prod\limits_{{\rm{}odd}\,i;i\not\in S}A^{+}_{x_{i}}\prod\limits_{{\rm even}\,j}A_{x_{j}=0}\right|\leq 1,\quad\left|\prod\limits_{{\rm{}odd}\,i;i\not\in S}A^{-}_{x_{i}}\prod\limits_{{\rm{}even}\,j}A_{x_{j}=1}\right|\leq 1.$ (35)
@@
$P(a_{2},\dots,a_{5}|x_{2},\dots,x_{5})=\textrm{PR}(a_{2},a_{3}|x_{2},x_{3})\textrm{PR}(a_{4},a_{5}|x_{4},x_{5}),$ (36)
straightforward evaluation leads to saturating the upper bounds, i.e., $A_{x_{2}=0}A^{+}_{x_{3}}A_{x_{4}=0}A^{+}_{x_{5}}=1$ and $A_{x_{2}=1}A^{-}_{x_{3}}A_{x_{4}=1}A^{-}_{x_{5}}=1$. Similar results hold for general $n\geq 3$.
@@
$I^{\frac{2}{n+1}}+J^{\frac{2}{n+1}}$$\leq$ $\prod\limits_{i\in S}(A^{+}_{x_{i}})^{\frac{2}{n+1}}+\prod\limits_{i\in S}(A^{-}_{x_{i}})^{\frac{2}{n+1}}$ (37)
@@
$\leq$ $\sin^{\frac{4|S|}{n+1}}\left(\frac{\pi}{2}-\frac{1}{|S|}\sum_{i\in S}\vartheta_{i}\right)+\sin^{\frac{4|S|}{n+1}}\left(\frac{1}{|S|}\sum_{i\in S}\vartheta_{i}\right)$ (38)
@@
where in Eq. (37) we have used that $|A^{+}_{x_{i}}|\leq\cos^{2}\vartheta_{i}$ and $|A^{-}_{x_{i}}|\leq\sin^{2}\vartheta_{i}$ for $\vartheta_{i}\in[0,\pi/2]$ as in the derivation of Eq. (14). The inequality (38) is from [31, Lemma 1], namely $\prod_{i=1}^{k}\sin^{\frac{1}{k}}\vartheta_{i}\leq\sin(k^{-1}\sum_{i=1}^{k}\vartheta_{i})$, where $k=|S|$. Note that, in this case, the relevant quantity for the bound is not the number of classical sources in the network, $\ell$, but the number of parties which only receive classical systems, $|S|$. In Eq. (39), we have defined $\vartheta=|S|^{-1}\sum_{i\in S}\vartheta_{i}$. The last inequality is from the fact that the unique extreme point of the continuous function $\cos^{\frac{4k}{n+1}}\vartheta+\sin^{\frac{4k}{n+1}}\vartheta$ for $\vartheta\in[0,\pi/2]$ is $\vartheta=\pi/4$.
@@
As in the case in the previous subsection, a chain network with an even number of parties can be decomposed into subnetworks consisting of $n-1$ parties. If there are $\ell$ classical sources, each subnetwork contain no less than $\ell-1$ classical sources, and thus the $n-1$-partite joint correlations satisfy the inequality (40) for $n-1$ parties.
@@ A.3 Quantum violations
For even $n$, consider two connected $n-1$-partite subnetworks $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, where $\mathcal{N}_{1}$ consists of $\textsf{A}_{1},\dots,\textsf{A}_{n-1}$ and $\mathcal{N}_{2}$ consists of $\textsf{A}_{2},\dots,\textsf{A}_{n}$. The strategy of [31] gives $\ell$-QNN correlations in each subnetwork, since it also gives the value in Eq. (41) for $\ell\geq n/2-1$.
@@ B.1 The tripartite chain
In quantum scenarios, consider the case where each source distributes a generalized EPR states given by Eq. (9). The total state is thus given by
$|\Phi\rangle=\cos\theta_{1}\cos\theta_{2}|0000\rangle+\cos\theta_{1}\sin\theta_{2}|0011\rangle+\sin\theta_{1}\cos\theta_{2}|1100\rangle+\sin\theta_{1}\sin\theta_{2}|1111\rangle.$
@@ B.2 Long chains
$=$ $\sum_{i_{2},\dots,i_{n}}\bigg{[}\left(\prod_{j=2}^{n-1}\alpha_{i_{j}}\right)|0,0i_{2},\dots,i_{n-1}0,0\rangle+\left(\prod_{j=2}^{n-1}\alpha_{\overline{i}_{j}}\right)|1,1i_{2},\dots,i_{n-1}1,1\rangle$
$\qquad\quad+\left(\prod_{j=2}^{n-1}\alpha_{i_{j}}\right)|0,0i_{2},\dots,i_{n-1}1,1\rangle+\left(\prod_{j=2}^{n-1}\alpha_{\overline{i}_{j}}\right)|1,1i_{2},\dots,i_{n-1}0,0\rangle\bigg{]},$
@@
$A^{(1)}_{0}=\cos\vartheta\sigma_{3}+\sin\vartheta\sigma_{1},\,A_{1}=\cos\vartheta\sigma_{3}-\sin\vartheta\sigma_{1}$
@@
$\text{CHSH}_{|\phi\rangle}:=$ $\,\left(A^{(1)}_{0}A^{(n)}_{2i}+A^{(1)}_{0}A^{(n)}_{2i+1}+A^{(1)}_{1}A^{(n)}_{2i}-A^{(1)}_{1}A^{(n)}_{2i+1}\right)_{|\phi\rangle}$ (46)

### v4

@@ VII Inequalities for general networks
$\left\langle\sum_{\vec{t}}M^{\vec{t}}\mathcal{S}^{(\vec{t})}\left(B_{x_{i_{1}}}^{(\vec{t})},\dots,B_{x_{i_{k}}}^{(\vec{t})}\right)\right\rangle\leq 2^{k-1}.$ (25)
VIII Noise robustness
@@
$\rho_{v_{i}}=v_{i}|\phi_{i}\rangle\langle{\phi_{i}}|+\frac{1-v_{i}}{d^{2}}I_{d^{2}},$
@@
Using the states $|\phi_{i}\rangle$ and the measurements that achieve the quantum violations described throughout the manuscript (see, e.g., Section IV and Appendix A.3), the corresponding quantities $I$ and $J$ get multiplied by a factor $\prod_{i=1}^{m}v_{i}$. This implies that, in order to observe a violation, the visibilities of the sources must satisfy $\prod_{i=1}^{m}v_{i}>\frac{1}{\sqrt{2}}$ when using maximally entangled states. Similarly, using maximally entangled states for $|\phi_{i}\rangle$ in the $n+1$-partite star network, $\ell$-NN correlations can be detected via the violation of Eq. (20) for $\prod_{i=1}^{m}v_{i}>2^{-\left(\ell-\frac{n}{2}\right)}$, while $\ell$-QNN correlations can be detected via the violation of Eq. (17) whenever $\prod_{i=1}^{m}v_{i}>2^{-\frac{\ell}{2}}$. As discussed in Section IV, this shows that Eq. (20) cannot be used for detecting FNN (although it can be used for detecting $\ell$-NN as long as $\ell>n/2$), while Eq. (17) can be used for detecting $\ell$-QNN in a noise-robust manner independently of the number of sources in the network, since the critical visibility only depends on $\ell$.
@@
$\prod_{i=1}^{n}v_{i}>\frac{1}{\sqrt{1+\prod_{j=1}^{n-1}\sin^{2}2\theta_{j}}}.$
Analogously, when decomposing star networks into sets of tripartite chain networks, FNN can in principle be guaranteed when every pair of visibilities satisfies
$v_{i}v_{j}>\frac{1}{\sqrt{1+\sin^{2}2\theta_{i}\sin^{2}2\theta_{j}}}.$
@@
IX Discussion
In this work we have provided Bell-like inequalities that witness genuinely quantum phenomena in networks, in the sense that their violations by particular distributions guarantee that these cannot be reproduced when at least a number of the sources in the network are classical. We have done so using two different assumptions on the nonclassical sources. First we assumed that they are only constrained by the no-signaling principle. This leads to a hierarchy of levels of network nonlocality, that interpolates from the standard definition of network nonlocality [23] to that of full network nonlocality [35]. Second, we assumed that the sources distribute systems described by quantum mechanics. This gives rise to a hierarchy analogous to the previous one, that has milder requirements for experimental observations and is thus more suitable for applications in quantum communication. More concretely, we have provided Bell-like inequalities whose violation guarantees (i) $n/2$-QNN and $\ell$-NN in the $n$-partite chain network (Eq. (14), and Eq. (30) in Appendix A, respectively), (ii) $\ell$-NN and $\ell$-QNN for any $\ell$ in the $n+1$-partite star network (Eqs. (20) and (17), respectively), and (iii) FNN in arbitrary chain and star networks (Eqs. (22) and (24), respectively).
Being able to guarantee, from observed statistics, the absence of classical sources in a network is an important step in the development of network-based quantum information protocols. Concrete protocols exist, for instance for quantum conference key agreement [49, 50, 51] or quantum secret sharing [52, 53]. However, their proofs of security rely on guaranteeing global multipartite nonclassicality, which leads to demanding requirements. Our work paves the way to the development of protocols and security proofs that exploit the network structure, which has received limited attention so far [54, 55, 53]. Indeed, consider the $n+1$-partite star network. Following Refs. [54, 53] one can bound the predictive power of an adversary for all outcomes $\vec{a}$ by the value obtained in Eq. (17) via
$D(P(\vec{e}|\vec{a};\vec{x},y,\vec{z}),\prod_{i=1}^{n}P(e_{i}|z_{i}))\leq(2-\mathcal{S}_{\ell})n,$
@@ IX Discussion
Having inequalities for detecting $\ell$-NN and $\ell$-QNN allows to easily see how inserting the assumption that nature is governed by quantum mechanics results in milder conditions to guarantee nonclassicality, which is relevant for experimental realizations. For instance, detection of FNN ($\ell=1$) in the star network with $n=3$ via Eq. (20) is not possible with noisy versions of the generalized EPR states considered throughout for any visibility, while detection of FQNN is possible when the product of the states’ visibilities exceeds $2^{-1/6}\approx 0.891$.
@@
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (Nos. 62172341, 61772437), Sichuan Natural Science Foundation (No. 2023NSFSC0447), Interdisciplinary Research of Southwest Jiaotong University China (No.2682022KJ004), the NCCR SwissMAP (funded by the Swiss National Science Foundation, grant no. 205607), the Spanish Ministry of Science and Innovation MCIN/AEI/10.13039/ 501100011033 (CEX2019-000904-S and PID2020-113523GB-I00), the Spanish Ministry of Economic Affairs and Digital Transformation (project QUANTUM ENIA, as part of the Recovery, Transformation and Resilience Plan, funded by EU program NextGenerationEU), Comunidad de Madrid (QUITEMAD-CM P2018/TCS-4342), Universidad Complutense de Madrid (FEI-EU-22-06), and the CSIC Quantum Technologies Platform PTI-001.
@@ Appendix A Nonlinear inequalities for chain networks
In this appendix we analyze the cases that are missing in Section III.2, namely, $\ell$-QNN in chain networks with an even number of parties, and $\ell$-NN in chain networks with an arbitrary number of parties. The core of the derivations is based on the concept of $k$-independent sets of sources [31].
@@ A.1 $\ell$-QNL bounds for even number of parties
Let us begin discussing the case when the nonclassical sources distribute quantum systems. An inequality for the case of chains with an odd number of parties was derived in Section III.2, namely
$I^{\frac{2}{n+1}}+J^{\frac{2}{n+1}}\leq\begin{cases}2^{\frac{n-1}{2(n+1)}}&\ell>\frac{n}{2}\\ \sqrt{2}&\ell<\frac{n}{2}\end{cases}.$
@@
If less than half of the sources in the network are classical, the situation is similar to that in Eq. (13). Namely, the joint correlations generated by either $\mathcal{N}_{1}$ or $\mathcal{N}_{2}$ achieve the maximal bound of $\sqrt{2}$ for the inequality (12) with $n-1$ parties, i.e., $I^{\frac{2}{n}}+J^{\frac{2}{n}}\leq\sqrt{2}$. This means that, in the case of an even number of parties and less than half of the sources being classical (i.e., for $\ell<n/2$), it is not currently possible to witness $\ell$-QNN.
@@ A.2 $\ell$-NN bounds
Now we focus on the bounds that are achieved when the only constraint on the nonclassical sources is that they distribute systems that satisfy the no-signaling principle. As in the previous section, we distinguish the cases of networks with odd and even number of parties.
@@
Suppose that the chain network consists of $\ell$ classical variables with $\ell\geq 1$ while the other sources distribute arbitrary no-signaling systems, as shown in Fig. 2(b). Equation (9) holds for this kind of networks. However, in contrast with the derivation in Section III.2, for sources distributing nonsignaling systems the objects in Eq. (10) are bounded by
$\left|\left\langle\prod\limits_{{\rm{}odd}\,i;i\not\in S}A^{+}_{x_{i}}\prod\limits_{{\rm even}\,j}A_{x_{j}=0}\right\rangle\right|\leq 1,$ (26)
$\left|\left\langle\prod\limits_{{\rm{}odd}\,i;i\not\in S}A^{-}_{x_{i}}\prod\limits_{{\rm{}even}\,j}A_{x_{j}=1}\right\rangle\right|\leq 1.$
@@
$P(a_{2},\dots,$ $a_{5}|x_{2},\dots,x_{5})$
$=\textrm{PR}(a_{2},a_{3}|x_{2},x_{3})\textrm{PR}(a_{4},a_{5}|x_{4},x_{5}),$
straightforward evaluation leads to saturating the upper bounds, i.e., $\left\langle A_{x_{2}=0}A^{+}_{x_{3}}A_{x_{4}=0}A^{+}_{x_{5}}\right\rangle=1$ and $\left\langle A_{x_{2}=1}A^{-}_{x_{3}}A_{x_{4}=1}A^{-}_{x_{5}}\right\rangle=1$. Similar results hold for general $n\geq 3$.
@@
$I^{\frac{2}{n+1}}+J^{\frac{2}{n+1}}$$\leq$ $\left\langle\prod\limits_{i\in S}A^{+}_{x_{i}}\right\rangle^{\frac{2}{n+1}}+\left\langle\prod\limits_{i\in S}A^{-}_{x_{i}}\right\rangle^{\frac{2}{n+1}}$ (28)
@@
$\leq$ $\sin^{\frac{4|S|}{n+1}}\left(\frac{\pi}{2}-\frac{1}{|S|}\sum_{i\in S}\vartheta_{i}\right)$
$+\sin^{\frac{4|S|}{n+1}}\left(\frac{1}{|S|}\sum_{i\in S}\vartheta_{i}\right)$
@@
where in Eq. (28) we have used that $\left|\left\langle A^{+}_{x_{i}}\right\rangle\right|\leq\cos^{2}\vartheta_{i}$ and $\left|\left\langle A^{-}_{x_{i}}\right\rangle\right|\leq\sin^{2}\vartheta_{i}$ for $\vartheta_{i}\in[0,\pi/2]$ as in the derivation of Eq. (11). The inequality (28) is from [31, Lemma 1], namely $\prod_{i=1}^{k}\sin^{\frac{1}{k}}\vartheta_{i}\leq\sin(k^{-1}\sum_{i=1}^{k}\vartheta_{i})$, where $k=|S|$. Note that, in this case, the relevant quantity for the bound is not the number of classical sources in the network, $\ell$, but the number of parties which only receive classical systems, $|S|$. In Eq. (29), we have defined $\vartheta=|S|^{-1}\sum_{i\in S}\vartheta_{i}$. The last inequality is from the fact that the unique extreme point of the continuous function $\cos^{\frac{4k}{n+1}}\vartheta+\sin^{\frac{4k}{n+1}}\vartheta$ for $\vartheta\in[0,\pi/2]$ is $\vartheta=\pi/4$.
@@
As in the case in the previous subsection, a chain network with an even number of parties can be decomposed into subnetworks consisting of $n-1$ parties. If there are $\ell$ classical sources, each subnetwork contains no less than $\ell-1$ classical sources, and thus the $n-1$-partite joint correlations satisfy the inequality (30) for $n-1$ parties.
@@ A.3 Quantum violations
For even $n$, consider two connected $n-1$-partite subnetworks $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, where $\mathcal{N}_{1}$ consists of $\textsf{A}_{1},\dots,\textsf{A}_{n-1}$ and $\mathcal{N}_{2}$ consists of $\textsf{A}_{2},\dots,\textsf{A}_{n}$. The strategy of Ref. [31] gives $\ell$-QNN correlations in each subnetwork, since it also gives the value in Eq. (31) for $\ell\geq n/2-1$.
@@ B.1 The tripartite chain
In quantum scenarios, consider the case where each source distributes generalized EPR states given by Eq. (8). The total state is thus given by
$|\Phi\rangle$ $=$ $\cos\theta_{1}\cos\theta_{2}|0000\rangle+\cos\theta_{1}\sin\theta_{2}|0011\rangle$ (32)
$+\sin\theta_{1}\cos\theta_{2}|1100\rangle+\sin\theta_{1}\sin\theta_{2}|1111\rangle.$
@@ B.2 Long chains
$=$ $\sum_{i_{2},\dots,i_{n}}\left[\left(\prod_{j=2}^{n-1}\alpha_{i_{j}}\right)|0,0i_{2},\dots,i_{n-1}0,0\rangle\right.$
$\qquad\quad+\left(\prod_{j=2}^{n-1}\alpha_{\overline{i}_{j}}\right)|1,1i_{2},\dots,i_{n-1}1,1\rangle$
$\qquad\quad+\left(\prod_{j=2}^{n-1}\alpha_{i_{j}}\right)|0,0i_{2},\dots,i_{n-1}1,1\rangle$
$\qquad\quad\left.+\left(\prod_{j=2}^{n-1}\alpha_{\overline{i}_{j}}\right)|1,1i_{2},\dots,i_{n-1}0,0\rangle\right],$
@@
$A^{(1)}_{0}=\cos\vartheta\sigma_{3}+\sin\vartheta\sigma_{1},\,A_{1}=\cos\vartheta\sigma_{3}-\sin\vartheta\sigma_{1},$
@@
$\text{CHSH}_{|\phi\rangle}\!:=$ $\left\langle\!A^{(1)}_{0}\!A^{(n)}_{2i}\!+\!A^{(1)}_{0}\!A^{(n)}_{2i+1}\!+\!A^{(1)}_{1}\!A^{(n)}_{2i}\!-\!A^{(1)}_{1}\!A^{(n)}_{2i+1}\!\right\rangle_{\!\phi}$ (37)
@@
$+2(p_{01}+p_{11})\min\limits_{\oplus_{i=1}^{n-1}a_{i}=01}\left\{\sqrt{1+r_{\vec{i}}^{2}\prod\limits_{j=1}^{n}\alpha^{2}_{i_{j}}}\right\}$